Ether, an organic compound that used to be used as an anesthetic, has a density of 0.736g/cm^3 at 20 degrees Celsius . What is the volume, in liters, of 225.0 grams of ether?
1
Expert's answer
2017-06-26T11:36:07-0400
m = 225.0 g
ρ = 0.736 g/cm^3
1 liter = 1000cm^3
V = m/ρ
V = 225 / 0.736 = 305.7 cm^3 = 0.3057 liters